# Seo Hyun Kim1     Kai Tzu-iunn Ong1 Taeyoon Kwon1     Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1

### v1

@@ 11 Acknowledgments
Appendix A Further Implementation Details
A.1 Cause-effect Commonsense Relations
@@
A.2 Algorithms for [Uncaptioned image]Theanine
@@
A.3 Computational Experiments
@@
Appendix B TeaFarm Evaluation
@@
Appendix C The [Uncaptioned image]TeaBag Dataset
We curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of:
@@ Appendix C The [Uncaptioned image]TeaBag Dataset
  • Two synthesized session 6 for each episode (thus 200 in total), each of which naturally leads to one of the counterfactual questions.
@@
Appendix D G-Eval
G-Eval (Liu et al., 2023) is a framework using LLMs with chain-of-thoughts (CoT) and a form-filling paradigm, to assess the quality of models’ text generation. G-Eval with GPT-4 has been shown to generate evaluation results that highly align with human judgement (Liu et al., 2023; Kim et al., 2024b). Due to the high API cost for it, we conduct head-to-head comparisons between Theanine and 8 baselines using 100 episodes from MSC and CC (around 3,600 Theanine’s responses being compared in each comparison with one baseline), rather than the whole test sets. The prompts for G-Eval are in Figure 2021, and 22. We use SciPy to calculate p-values. [^10] https://scipy.org/
Appendix E Dataset-Specific G-Eval Results
We provide dataset-specific results of G-Eval in Table 4 and 5.
Appendix F Human Evaluation
@@
  • Appropriateness of relation-aware memory linking: In this evaluation, we ask the workers to judge whether they agree that the relation-aware linking is properly done for two given memories.
  • Helpfulness of context-aware timeline refinement: This evaluation requires the workers to determine if they agree that our context-aware refinement really tailors a raw timeline into a resource of useful information for generating the next response.
  • The quality of Theanine’s responses: Here, the workers are asked to judge if the responses of Theanine correctly refer to past conversations. After reading our responses and past memories, they choose whether the responses entail, contradict, or are neutral to past memories. To improve evaluation quality, we use GPT-4 to select responses for this specific evaluation based on past memories, addressing the fact that not every turn in the conversation requires previous information to generate the next response (In the other two evaluations, the samples are randomly selected).
Each data sample is judged by 3 different workers, and we report the results based on the majority rule. In the third evaluation, when every option (entailment, neutral, contradiction) gets one vote, we consider it neutral (13 samples in total). These human evaluations are conducted on 300 samples. The interfaces provided to AMT workers, which include detailed instructions for human evaluation, are shown in Figure 910, and 11.
@@ Appendix G Empirical Examples
We provide several empirical examples of [Uncaptioned image]Theanine. Examples of relation-aware memory linking are in Figure 1213, and 14. Examples of utilizing refined timeline for response generation are in Figure 15.
@@ Appendix H Prompts
  • + Memory update (baseline): Figure 19.
@@
  • G-Eval: Figure 20,  21, and  22.
@@
Appendix I Terms for Use of Artifacts
@@
Datasets: Multi-session Chat (MSC)
Dialogue Session: Session 3 Session 4 Session 5 Avg.
[Uncaptioned image]Theanine Theanine’s Winning Percentage
          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
All Dialogue History 74% 79% 72% 78% 80% 74% 74% 76% 71% 75% 78% 72%
All Memories and $\mathcal{D}$ 71% 66% 64% 72% 69% 67% 70% 65% 63% 72% 66% 65%
     + Memory Update 59% 54%* 58% 58% 53% 58% 56%* 54%* 57%* 58% 54%* 58%
Memory Retrieval 68% 61% 65% 73% 67% 67% 72% 64% 65% 71% 64% 66%
     + Memory Update 56%* 53% 57% 61% 58% 61% 53%* 53% 53% 57% 55% 57%
RSum-LLM 44% 57% 68% 41% 54%* 67% 43% 53% 63% 42% 55% 66%
MemoChat 65% 75% 70% 64% 69% 72% 65% 74% 72% 65% 73% 71%
COMEDY 72% 64% 84% 72% 61% 84% 69% 60% 81% 71% 61% 83%
Table 4: Winning percentage of Theanine in head-to-head G-Eval (of all generated responses) with baselines. Tested with Multi-Session Chat (MSC). Color gray shows where ours loses more. Underline: p-value $<0.001$. *: $0.001<$ p-value $<0.05$.
Datasets: Conversation Chronicles (CC)
Dialogue Session: Session 3 Session 4 Session 5 Avg.
[Uncaptioned image]Theanine Theanine’s Winning Percentage
          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
All Dialogue History 69% 85% 83% 67% 82% 79% 71% 82% 82% 69% 83% 81%
All Memories and $\mathcal{D}$ 66% 62% 68% 66% 69% 70% 69% 71% 73% 67% 68% 70%
     + Memory Update 54%* 59% 62% 55%* 58% 59% 55%* 63% 64% 54% 60% 62%
Memory Retrieval 67% 64% 71% 70% 67% 68% 71% 66% 71% 69% 66% 70%
     + Memory Update 57% 59% 61% 55%* 57% 62% 57% 62% 66% 56% 60% 63%
RSum-LLM 31% 64% 69% 29% 60% 67% 32% 60% 68% 31% 61% 68%
MemoChat 56%* 73% 73% 57% 69% 71% 55%* 70% 70% 56% 70% 71%
COMEDY 59% 44% 73% 57% 46% 71% 55%* 46% 71% 57% 46% 72%
Table 5: Winning percentage of Theanine in head-to-head G-Eval (of all generated responses) with baselines. Tested with Conversation Chronicles (CC). Color gray shows where ours loses more. Underline: p-value $<0.001$. *: $0.001<$ p-value $<0.05$.
@@
Figure 20: The prompt for the G-Eval: Specificity.
Figure 21: The prompt for the G-Eval: Memorability.
Figure 22: The prompt for the G-Eval: Consistency.

### v2

@@ 11 Acknowledgments
Appendix A Appendix Contents
  • Appendix B.1: Cause-effect Commonsense Relations.
  • Appendix B.2: Algorithms for [Uncaptioned image]Theanine.
  • Appendix B.3: Implementation Details on Computational Experiments.
  • Appendix C: TeaFarm Evaluation.
  • Appendix D: The [Uncaptioned image]TeaBag Dataset.
  • Appendix E Details on Evaluation Scheme 1 (G-Eval and Human Evaluations).
  • Appendix F: Session-specific Evaluation Results.
  • Appendix G: Empirical Examples.
  • Appendix H: Prompts Used in This Work.
  • Appendix I: Further Analyses and Discussions.
  • Appendix J: Terms for Use of Artifacts.
Appendix B Further Implementation Details
B.1 Cause-effect Commonsense Relations
@@
B.2 Algorithms for [Uncaptioned image]Theanine
@@
B.3 Implementation Details on Computational Experiments
@@
Appendix C TeaFarm Evaluation
@@
Appendix D The [Uncaptioned image]TeaBag Dataset
As a byproduct of TeaFarm, we curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of:
@@ Appendix D The [Uncaptioned image]TeaBag Dataset
  • Two synthesized follow-up conversations (i.e., session 6) for each episode (thus 200 in total), each of which naturally guides the conversation from session 5 towards one of the counterfactual questions.
@@
For the overview of data collection, please refer to step 1-4 of TeaFarm (Figure 10).
Appendix E Details on Evaluation Scheme 1
We perform evaluations using sessions 3-5 from MSC and CC, as all settings are almost identical before the end of session 2, due to the fact that there is no memory to update before then.
The test sets of MSC and CC contain over 500 and 20,000 episodes of conversations, where each episode has 5 dialogue sessions, yielding 1.2M turns of responses in total. Due to the limited budget for generation (both baselines and ours), when not specified, we sample 50 episodes from each dataset for experiments in this paper (around 3.6K conversational turns in total).
E.1 G-Eval
G-Eval (Liu et al., 2023) is a framework using LLMs with chain-of-thoughts (CoT) and a form-filling paradigm to assess the quality of models’ text generation. G-Eval with GPT-4 has been shown to generate evaluation results that highly align with human judgement (Liu et al., 2023; Kim et al., 2024b) and thus has been widely applied in many LM-based projects. We conduct G-Eval on 5 episodes.
The prompt for evaluating the helpfulness of retrieved memories is in Figure 23. We use SciPy to calculate p-values. [^12] https://scipy.org/
E.2 Human Evaluation
@@
  • Appropriateness of relation-aware memory linking: In this evaluation, we ask the workers to judge whether they agree that the relation-aware linking is properly done for two given memories. The interface provided to AMT workers, which includes detailed instructions for human evaluation, is shown in Figure 11.
  • Helpfulness of context-aware timeline refinement: This evaluation requires the workers to determine if they agree that our context-aware refinement really tailors a raw timeline into a resource of useful information for generating the next response. The interface provided to AMT workers, which includes detailed instructions for human evaluation, is shown in Figure 12.
  • The quality of responses: Here, the workers are asked to judge if the responses correctly refer to past conversations. After reading our responses and past memories, they choose whether the responses entail, contradict, or are neutral to past memories. To improve evaluation quality, we use GPT-4 to select responses for this specific evaluation based on past memories, addressing the fact that not every turn in the conversation requires previous information to generate the next response (In the other two evaluations, the samples are randomly selected). The interface provided to AMT workers, which includes detailed instructions for human evaluation is shown in Figure 13.
  • The helpfulness of retrieved memories: Given the same dialogue context, human workers are asked to select a memory that is more helpful for generating a next response from ours’ and a baseline’s retrieval. The interface provided to AMT workers, which includes detailed instructions for human evaluation is shown in Figure 14
Each data sample is judged by 3 different workers, and we report the results based on the majority rule. In the third evaluation, when every option (entailment, neutral, contradiction) gets one vote, we consider it neutral (13 samples in total). These human evaluations are conducted on 100 conversational turns.
Appendix F Session-specific Evaluation Results
We provide session-specific results for automatic evaluations in Table 5.
@@ Appendix G Empirical Examples
We provide several empirical examples of [Uncaptioned image]Theanine. Examples of relation-aware memory linking are in Figure 1516, and 17. Examples of utilizing refined timeline for response generation are in Figure 18.
@@ Appendix H Prompts
  • Memory Update (baseline): Figure 22.
@@
  • G-Eval: The prompt for evaluating the helpfulness of retrieved memories is in Figure 23.
@@
Appendix I Further Analyses and Discussions
Memory Summarization.
At the end of each session, we use ChatGPT (gpt-3.5-turbo-0125) to summarize the conversation into memory sentences. We conduct examinations on such conversation-memory summarization using 100 randomly sampled sessions from MSC and CC to make sure the quality of raw memories is acceptable. The result is in Table 4.
Memories that … No Can’t judge Yes
Contain faulty statements 90% 9% 1%
Miss important statements 95% 4% 1%
Table 4: Human evaluation of conversation-to-memory summarization in Theanine.
Appendix J Terms for Use of Artifacts
@@ Appendix I Further Analyses and Discussions
Datasets: Multi-session Chat (MSC) & Conversation Chronicals (CC)
Session: Session 3 Session 4 Session 5
Methods / Metrics B-4 R-L Mauve Bert B-4 R-L Mauve Bert B-4 R-L Mauve Bert
All Dialogue History 3.13 18.04 17.34 87.17 3.17 17.96 18.54 87.12 3.53 18.69 17.42 87.31
All Memories & Current Context $\mathcal{D}$ 2.69 17.29 28.30 87.10 3.10 17.38 22.52 87.06 3.16 17.75 22.35 87.21
+ Memory Update (Bae et al., 2022) 2.80 17.51 22.92 87.11 2.88 17.24 21.22 86.99 3.16 17.90 22.04 87.24
Memory Retrieval (Xu et al., 2022a) 3.44 18.33 24.68 87.30 3.38 17.55 21.95 87.17 3.46 18.31 19.70 87.33
+ Memory Update (Bae et al., 2022) 3.10 18.08 25.02 87.24 2.99 17.37 25.97 87.10 3.11 17.78 20.99 87.28
Rsum-LLM (Wang et al., 2023) 0.83 11.30 2.45 85.25 0.87 11.35 2.32 82.20 0.90 11.78 2.33 85.30
MemoChat (Lu et al., 2023) 1.88 14.83 14.56 86.57 1.81 14.27 10.57 86.43 1.91 14.96 9.13 86.56
COMEDY (Chen et al., 2024b) 1.14 12.80 4.74 85.53 1.57 13.18 5.16 85.56 1.42 13.56 3.94 85.69
[Uncaptioned image]Theanine (Ours) 4.21 19.21 45.53 87.70 4.42 18.63 37.84 87.52 4.34 19.23 41.18 87.70
Table 5: Session-specific results of agent performance in response generation.
@@
Figure 14: Interface for human evaluation regarding the helpfulness of retrieved memories.
@@
Figure 23: The prompt for the G-Eval: Helpfulness of Retrieved Memories.